40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 167 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  9196   Thu Oct 3 09:49:49 2013 SteveUpdateSUSETMY oplev laser is sick

Quote:

As another proof that sometime is ill with ETMX Optical Lever:

We scanned the ETMX bias in PIT using ezcastep and saw that the OL response is very screwy. In the attached, you can see that the ETMX SUSPIT signal shows that the actual motion is good and linear. In fact, our sus diagonalization is extermely good and there's almost no signal in SUSYAW.

 ETMY oplev laser clearly showing a tail when it was projected up the sealing.

PS (10-4-2013): I checked the beam quality again as it was removed from the table: it had a good image at 3 meters

  9195   Thu Oct 3 09:01:06 2013 manasaUpdateGreen LockingALS high frequency noise

 As I was trying to solve the 2 arm ALS problem, I found the Y arm ALS not so stable AGAIN :( . I measured the in-loop noise of the X arm as ~400Hz/rtHz (60 picometers).

I went ahead and checked the out of loop noise of the ALS and found there is some high frequency noise creeping in above 20Hz for the Y arm ALS (blue curve). I checked the UGFs and phase margins of the phase tracker loops and found they were good (UGF above 1.4KHz and phase margins between 40 and 60 degrees).

So the suspect now is the PDH servo loop of both the arms which has to be checked.

Attached is the out-of loop noise plots of X and Y arm ALS.

Attachment 1: ALS_outloop.pdf
ALS_outloop.pdf
  9194   Thu Oct 3 08:57:00 2013 jamieUpdateComputer Scripts / Programspianosa can't find Jamie PPA

Quote:

Message on 'pianosa':

Failed to fetch http://ppa.launchpad.net/drgraefy/nds2-client/ubuntu/dists/lucid/main/binary-amd64/Packages.gz  404  Not Found

Sorry, that was an experiment to see if I could set up a general-use repository for the NDS packages.  I've removed it, and did an update/upgrade.

  9193   Thu Oct 3 02:51:26 2013 JenneUpdateLSCPRMI locked "forever", some ALS fiddling

First up for me this evening was getting the PRMI locked.

I used the IFO configure screen to lock the X and Y arms, then aligned them using the ASS scripts.  Then used the IFO config screen to restore the Michelson, and did some fine tune tweaking of the BS alignment by looking at the AS camera.  Then, I restored the PRMI from the IFO config screen, tweaked the PRM a little bit in yaw, and was able to get a lock using REFL 165 I&Q for ~25 minutes before I got bored and unlocked things.  I used the ASS for the PRM to align the PRM, then turned off the ASS.  POP110 and POP22 both drifted down, but by a small amount, and at the end (when I turned the ASS back on for PRM), they picked back up to about their original levels.

PRMI_lock_ASSatBeginningAndEnd.png

(Note to self: to get it to print both plots, chose custom paper size, make it 14.5 by 11.  Don't ask why, just do it, because it works.  Also, in PNG device properties, increase the compression to 9.)

 

After I played with the PRMI, I started looking at the ALS system. 

I had both arms locked on IR using the regular LSC system (so POX and POY for the error signals).  Then I opened up the green shutters, and got both arms locked on green (so the green lasers were just following the arms...no digital ALS business).  I went out to the PSL table and tweaked up the alignment of the green beams (didn't need much at all, just an itsy bitsy bit in yaw, mostly).  I saw a very strong peak for the Yarm vs. PSL (around -19dBm), and there was a harmonic of that beat.  Opening and closing the Xarm green shutter had no effect on these peaks, so there wasn't any kind of X-Y cross beat sneaking around that I could see.  That's really as far as I got - I think (but haven't checked) that Manasa may have removed the power splitter / combiner, so that the RF analyzer is only looking at the Y beat PD (she mentioned earlier today that she was going to give that a try to narrow things down). 

After that, Rana and I went back to the PRMI for some noise stuff, and worked on the PMC.  See those separate elogs for info on those activites.

  9192   Thu Oct 3 02:46:58 2013 ranaMetaphysicsPEMUSGS Furlough

ObamaCare_GOP.png

  9191   Thu Oct 3 02:43:34 2013 rana, jenneSummaryLSCPRMI: comparison of 1f and 3f signals w/ calibration

The attached plot shows the spectra of all the REFL signals with the PRMI SB lock.

We excited the ITMY_LSC with 3000 counts. We used the Masayuki calibration of ITMY (5 nm / count * (1/f^2)) to estimate this peak in the REFL spectra.

To correctly scale the REFL spectra we account for the fact that the DTT BW was "0.187 Hz" and we turn off the "Bin" radio box before measuring the peak height with the cursor.

Since the ITMY motion is 3000 * 5e-9 / (580.1 Hz)^2 = 44.6 pm_peak, we want the DTT spectrum of the REFL spectra to report that too.

i.e. to convert from peak height to meters_peak, we use this formula:

meters_peak = peak_height * sqrt(BW) * sqrt(2)

I *think* that since the line shows up in multiple bins of the PSD, we should probably integrate a ~0.5 Hz band around the peak, but not sure. Need to check calibration by examining the time series, but this is pretty close.

Mystery: why are the REFL_I 3f signals nearly as good in SNR as the 1f signals? The modelling shows that the optical gain should be ~30-100x less. Can it be that our 1f electronics are that bad?

Bonus: notice how we have cleverly used the comb of bounce frequencies around the calibration line to determine that REFL11 is clipping!

Attachment 1: REFL_signals_CalLinesLinedUp.pdf
REFL_signals_CalLinesLinedUp.pdf
  9190   Thu Oct 3 01:24:31 2013 Jenne, RanaUpdateIOOPMC

The PMC transmission was around 0.78 all day, rather than the usual 0.83ish.  Rana went out to the PSL table and fixed up the PMC alignment.  This should not need to be done very often, so things to check before touching the alignment are FSS / PMC settings (digital stuff).  Make sure that the PC RMS (on the FSS screen) is low (at least below 2, preferably below 1), and that the FSS Fast monitor is near 5ish (not near 0 or 10).  

This is a capture of PMC REFL's camera after Rana was finished. If it doesn't look this good when you finish then you are not done. Never do PMC alignment without looking at the PMC REFL camera.

PMCR_1064822387.bmp

The attached trend shows 80 days of PMC REFL and TRANS. The bad alignment stuff started on Sep 21-24 time period. You know who you are.

Attachment 2: pmc80.png
pmc80.png
  9189   Thu Oct 3 01:18:57 2013 KojiHowToLSCsteps to full IFO

I vote on PRMI+1arm -> PRFPMI

  9188   Thu Oct 3 01:06:48 2013 rana, jenneUpdateSUSoplev XY-plots reflect new calibration

As another proof that sometime is ill with ETMX Optical Lever:

We scanned the ETMX bias in PIT using ezcastep and saw that the OL response is very screwy. In the attached, you can see that the ETMX SUSPIT signal shows that the actual motion is good and linear. In fact, our sus diagonalization is extermely good and there's almost no signal in SUSYAW.

Attachment 1: etmx_ol.png
etmx_ol.png
  9187   Thu Oct 3 00:01:59 2013 rana, jenneHowToLSCsteps to full IFO

In moving now to full IFO locking, there are a number of sub-states to diagnose:

  1. PRMI + 1 arm
  • Measure sensing matrix as arm is scanned into resonance. Compare time series of sensing matrix elements with New LoopTickle simulation. But first, we need more than 1 LOCKIN screen in the LSC! That will allow us to measure all of the elements of front_matrix.jpg simulataneously.
  • Measure 3f PRMI noise spectra as a function of arm position. Look for trouble.
  1. DRMI + 1 arm
  • Same as PRMI above.
  • Want to find why this is unstable sometimes. Make stable for t > 10 minutes.
  • Maybe add some QPD->ASC for SRC angular control, but how? Will this still work after the arms are resonant or will it be swamped by carrier contrast defect? Will Berlusconi ruin all of the Italian gelateria? Only time can tell...
  1. FPMI (non optically recombined) for ALS diagnosis
  2. PRFPMI (iLIGO configuration)
  • this ought to be easier than DRFPMI
  • will let us tell if our ALS is good enough to handle the coupled cavity pole
  1. DRFPMI (aLIGO style)

Which to do first and in what order?

  9186   Wed Oct 2 23:21:54 2013 ranaUpdateComputer Scripts / Programspianosa can't find Jamie PPA

Message on 'pianosa':

Failed to fetch http://ppa.launchpad.net/drgraefy/nds2-client/ubuntu/dists/lucid/main/binary-amd64/Packages.gz  404  Not Found

  9185   Wed Oct 2 14:25:39 2013 SteveUpdateVACvertex crane is still out of order

 

The smoke alarms were turned off and surrounding areas were covered with plastic.

The folding I-beam was ground down to be in level with the main beam.

Load bearing cable moved into correct position. New folding spring installed.

Crane calibration was done at 500 lbs at the end of the fully extended jib.

Than we realized that the rotating wheel limit switch stopped working.

This means that the crane is still out of order.

 

Attachment 1: grindingIbeam.jpg
grindingIbeam.jpg
Attachment 2: cableMoved.jpg
cableMoved.jpg
Attachment 3: 500lbs@end.jpg
500lbs@end.jpg
  9184   Tue Oct 1 19:42:19 2013 ranaSummaryCDSmegatron upgrade

Max and I started upgrading megatron to Ubuntu 12.NN today. We were having some troubles with getting latest python code to run to support the Summary pages stuff.

Its also a nice test to see what CDS tools fail on there, before we upgrade the workstations to Ubuntu 12.

Since its Linux, none of the usual upgrading commands worked, but after an hour or so of reading forums we were able to delete some packages and all the 3rd party packages and get the upgrade to go ahead. We'll have to re-install the LSC, GDS, LAL repos to get it back into shape and get NDS2 working. The upgrade is running in a 'screen' command on there.


Wed Oct 02 14:50:16 2013 

Update #1: The upgrade asks a couple dozen questions so it doesn't proceed by itself. I've been checking in to the 'screen' every couple hours to type in 'Yes' to let it keep going.


Update #2: It finished a few hours ago:

controls@megatron:~ 0$ uname -a
Linux megatron 3.2.0-54-generic #82-Ubuntu SMP Tue Sep 10 20:08:42 UTC 2013 x86_64 x86_64 x86_64 GNU/Linux
controls@megatron:~ 0$ date
Wed Oct  2 18:33:41 PDT 2013

  9183   Tue Oct 1 17:14:53 2013 masayukiUpdateGreen LockingALS servo filters modified

 

 [Manasa, Masayuki]

[revised at 10/1 pm 5:00]

As we mentioned in previous entry (elog#9171), the phase margin of ALS control was at most 20 degree. We modified the filter of C1ALS_XARM and C1ALS_YARM. The OLTF is in attachment1. Now the phase margins of both arms are more than 35 degree. I modified the FM5 filters of both servo.

FM5 filter is the filter for the phase compensation. It had the one pole at 1000 Hz and one zero at 1Hz. As you can see in attachment2, it start to lose the phase at 50 Hz. But the UGF of our ALS control loop is higher than 100 Hz, so I changed the pole from 1 kHz to 3 kHz in order to get more phase margin at UGF. The new servo have 10dB larger gain than previous filter at higer than 1kHz, but the control loop do nothing in that region, so it's no problem.

We have phase lag between 2 arms. I used same filters for both arms, so I'm wondering where these phase lag came from.

 

Attachment 1: OLTF.pdf
OLTF.pdf
Attachment 2: filter_change.pdf
filter_change.pdf
  9182   Tue Oct 1 14:12:22 2013 ranaSummaryCDSsvndumpfilter on linux1 makes NFS slow

 Yesterday and this morning's slow NFS disk access was caused by 'svndumpfilter' being run at linux1 to carve out the Noise Budget directory. It is being moved to another server; I think the disk access is back to normal speed now.

  9181   Tue Oct 1 11:24:02 2013 ranaUpdatePEMparticle counts

 Probably more important is to establish quantitatively how particle counts affects the lock acquisition or noise in the interferometer.

We don't want to adopt a "Sky is Falling" mentality as was done previously in LIGO when people were trying to outlaw burritos and perfume.

Dust monitor counts and human noses do not correlate well with the interferometer's nose.

pepelepew.jpg

When the vacuum system is closed, we might check to see if particle counts correlate with losses in the PMC or excess scatter on the ISC tables. If not, we should move on to other concerns.

  9180   Tue Oct 1 10:06:27 2013 SteveUpdatePEMparticle counts

How can we improve the cleanliness of the 40m IFO room 104 ? May be Mario Batali crocs if you chief chef of the lab likes it. We can do more effective things.

Atm 1 is showing the lab air quality dependence on outside air  measured at the top of the IOO chamber. This made me to measure the   differential  pressure between lab and out side.

Properly sized air conditioner is over pressurizing it's room by 0.05 [" Water] like clean assembly room.

The 40m at Vertex location is barely getting pressurized to 0.01 " W

The control room is OK with 0.04 " W but it's air quality is very bad! It is 10x worse at 0.5 micron than IFO room.

Every entry from well pressurized control room into barely pressurized  IFO  pumps dirty air to our "clean room."

This door should be closed.

The drill- room entry is ideal because it is using the same air conditioner as the  main lab, therefore it has cleaner air.

 

Things to do: seal holes of CES walls, seal paint wall at south end so it will shed less, replace gas kits on doors.

I have plugged the cable tray space entering the control room above Rosalba.

Attachment 1: 3xJuly4Cal.png
3xJuly4Cal.png
Attachment 2: whichOne.jpg
whichOne.jpg
  9179   Tue Oct 1 09:51:10 2013 ranaUpdateGreen LockingALS autolocker flowchart

  I think we can use the IMC autolocker to start with getting this started. Once Jamie fixes the NDSSERVER environment variable bug, we should be able to use his more slick automation code to make it auto lock.

  9178   Mon Sep 30 23:56:19 2013 manasaUpdateGreen LockingALS autolocker flowchart

[Masayuki, Manasa] 

Flowchart for ALS autolocker. The error signal thresholds will be decided by trial and error.

 ALSautolocker.png

  9176   Mon Sep 30 17:55:45 2013 manasaUpdateGreen LockingX and Y arm transmission needs to be decoupled

[Masayuki, Manasa]

Problem
We wanted to lock both the arms using ALS and get IR to resonate while arms are held using ALS. The X arm was locked using ALS and offsetter2 was used to scan the arm and find IR resonance. The Y arm was locked using ALS. But as the Y arm was brought closer to IR resonance, the X arm ALS loses lock. (attachment 1)

Discussion
We believe that this comes from the X and Y transmission not being well separated at the PSL table. The PBS is not sufficient to decouple them (A strong beatnote ~35dB between the X and the Y arm green lasers can be seen on the spectrum analyzer).

Solution
Decouple the X and Y arm transmitted beams at the PSL table. I am trying to find a wedged mirror/window that can separate the 2 beams at the PSL table before the beat PD (sadly the laseroptik HR532nm optics have no wedge)

 

Attachment 1: scan2.png
scan2.png
  9175   Mon Sep 30 13:02:51 2013 Masayuki,ManasaUpdateIOOPMC and MC alignment

[Manasa, Masayuki]

The MC lost lock around 8+hrs ago. The transmission from PMC was 0.77 this morning, so we aligned the PSL to the PMC using the two steering mirrors before the PMC. We brought the PMC transmission to 0.841. We also aligned the MC, and the MC transmission reflection now is 0.59.

  9174   Mon Sep 30 11:33:15 2013 ranaUpdateLSCLSC calibration screen

Quote:

  I fixed the XARM and YARM real time calibration servo.

I also change the C1CAL_MICH_A servo. Now the actuator response and the suspension TF are combined together and that filter name is BS_act. C1CAL_XARM_A and C1CAL_YARM_A have same kind of filters, ETMX_act and ETMY_act.

There are AI filter in each A servo and inv_AA, inv_DAA filters in CINV servo, but it's doesn't work correctly yet.

 These aren't servos. What he means is that he's changed some filters in the real time calibration screens so as to make the actuation and sensing parts more accurate, but the inversion of the AA filters is not accurate yet.

  9173   Mon Sep 30 08:53:22 2013 SteveUpdateVACvertex crane repair schedule has changed

Quote:

Quote:

 The folding beam removed as shown. Two man supporting it while I hammering it out. Pin was dry and it gulled into supporting hinges.

The rotating hinge will be machined and bushing will be added with Zerk fitting or similar. This will allow lubrication in the future.

 see elog #9111

Atm1,  The folding arm is back on with 0.1" misalignment at no load in the trolly's way. The other side of the I beam is 0.02" higher than the main beam.. New bushing and pin were greased up  with Krytox before installation.

The axial Zerk 1/8" pipe in the pin upper end can not take any fitting. There is no room. It is taped off.

This gap comes down to ~ 1/16" at fully extended arm with 225 lbs load at the end of it.

The present plan is to grind down the  the misalignment of 0.1"  for a slow-loaded trolly.

Steve Baker of Konecranes will be back to grind down this ridge and load test at 500 lbs on Tuesday,  OCT  1, 2013

 

 

 

 Konecranes rescheduled the completion of the Vertex crane to Wednesday, Oct. 2

  9172   Fri Sep 27 21:01:50 2013 MasayukiUpdateLSCLSC calibration screen

 

 I fixed the XARM and YARM real time calibration servo.

I also change the C1CAL_MICH_A servo. Now the actuator response and the suspension TF are combined together and that filter name is BS_act. C1CAL_XARM_A and C1CAL_YARM_A have same kind of filters, ETMX_act and ETMY_act.

There are AI filter in each A servo and inv_AA, inv_DAA filters in CINV servo, but it's doesn't work correctly yet.

  9171   Fri Sep 27 20:28:10 2013 manasaUpdateGreen LockingALS servo

[Masayuki, Manasa]

I. ALS servo loops
After fixing things with the phase tracking loop, we checked if things were good with the ALS servo loops.
We measured the OLTF of the X and Y arm ALS servo loops. In both cases the phase margin was ~20 degrees. There was no room to set enough phase margin. So we looked at the servo filters. We tried to modify the filters so that we could bring enough phase margin, but could not get at it. So we put back the old filters as they were.

 attachment1: OLTF of the ALS XARM and YARM control loops

attachment2: Current phase budget. FM4 and FM10 are the boost filters.

II. ALS in-loop noise
Also, I found that the overall noise of the ALS servo has gone up by about two orders of magnitude (in Hz/rtHz) over the whole range of frequencies for both the arms from the last time the measurements were made. I suspect this could be from some change in the calibration factor. Did anybody touch things around that could have caused this? Or can somebody recollect any changes that I made in the past which might have affected the calibration? Anyways, I will do the calibration again.

 

 

Attachment 1: OLTF.pdf
OLTF.pdf
Attachment 2: phase_badget_xarm_ALS.pdf
phase_badget_xarm_ALS.pdf
  9170   Fri Sep 27 16:02:23 2013 manasaUpdateGreen LockingY arm ALS phase tracking loop gain changed

[Masayuki, Manasa]

While trying to lock the arms using ALS we found that the locks were not very stable and the in-loop noise was higher than seen before.

I looked into things and checked the out-of loop noise for ALS and found that the Y arm ALS noise (rms) was higher than the X arm.

To troubleshoot, I measured the OLTF of the phase tracking loop. While X arm was healthy, things weren't looking good for the Y arm. Sadly, the Y phase tracking loop gain was set too high with a phase margin of -2 degrees. We brought down the gain from 300 to 150 and set the phase margin close to ~55 degrees.

X arm Phase tracker loop:
UGF = 1.8 K Hz
Phase margin = 50 degrees

Y arm Phase tracker loop:
UGF = 1.6 KHz
Phase margin = 55 degrees

Attachment 1: outofloop.pdf
outofloop.pdf
Attachment 2: PTX_OLTF.pdf
PTX_OLTF.pdf
Attachment 3: YPT_OLTF_after1.pdf
YPT_OLTF_after1.pdf
Attachment 4: YPT_OLTF_before.pdf
YPT_OLTF_before.pdf
  9169   Fri Sep 27 13:55:11 2013 MasayukiUpdateLSCLSC calibration screen

  I added the DAQ channel to all output of calibration servo. The name of channels are C1CAL_(plant name)_W_OUT_DQ.

I recompiled and restarted the model. Also I committed the changes to the svn of the calibration model.

  9168   Fri Sep 27 00:48:53 2013 MasayukiUpdateLSCFPMI noise caused by ARM locking

Quote:

Hidden in Nakano-kun's previous entries was that the phase margin of the X-Arm was only 9 degrees!! This extremely close to instability and makes for huge gain peaking. The feedback loop is increasing noise above 100 Hz rather than suppress. After some tweaks of the LSC filters we got a much more stable loop/.

So we today started to examine the sources of phase lag in the arm cavity sweeps. There were a few unfortunate choices in the XARM LSC filter bank which we tuned to get less delay.

Then I wrote a bunch of detail about how that worked, but the ELOG ate my entry because it couldn't handle converting my error signal noise plot into a thumbnail. Then it crashed and I restarted it. We also have now propagated the changes to the Y arm by copy/paste the filters and the result there is pretty much the same: low phase margin is now 38 deg phase margin. Noise is less bad.

 

[Rana, Masayuki

 I made the plot of the phase of the digital filters which Rana change and also  of the AA, AI, DAA, DAI filters. Now the biggest phase delay come from the timedelay of the digital system.

phase_badget.png

The UGF is around 150 Hz at that frequency the time delay has biggest phase  delay. Second one is the FM9 filter (this filter is BOOST filter). Then we have the AA filter, AI filter and so on, but these delay is roughly 5 degree.

As I said in previous entry, the time delay of the XARM control is roughly 300 usec, and we have 120 usec even only in C1SUS. Also between the C!SUS and C1LSC we have another 120 usec time delay. We want to increase the UGF to 300 Hz but because of the time delay of the digital system we cannot increase. So we should fix this problem.

 

After changing these filters, the FPMI noise is become better at high frequency. Before we have peak around the 100 Hz (because of 8 degree phase margin...), but they are gone. i attached the noise spectrum. This plot is measured by the real time calibration output. But even then, you can see the extra noise around 100 Hz in FPMI conpare to only MICH.

 FPMI_noises.png

 

 

  9167   Thu Sep 26 23:02:40 2013 ranaUpdateLSCFPMI noise caused by ARM locking

Hidden in Nakano-kun's previous entries was that the phase margin of the X-Arm was only 9 degrees!! This extremely close to instability and makes for huge gain peaking. The feedback loop is increasing noise above 100 Hz rather than suppress. After some tweaks of the LSC filters we got a much more stable loop/.

So we today started to examine the sources of phase lag in the arm cavity sweeps. There were a few unfortunate choices in the XARM LSC filter bank which we tuned to get less delay.

Then I wrote a bunch of detail about how that worked, but the ELOG ate my entry because it couldn't handle converting my error signal noise plot into a thumbnail. Then it crashed and I restarted it. We also have now propagated the changes to the Y arm by copy/paste the filters and the result there is pretty much the same: low phase margin is now 38 deg phase margin. Noise is less bad.

 

Attachment 1: Xarm_sweep_130926.pdf
Xarm_sweep_130926.pdf
Attachment 2: lsc.pdf
lsc.pdf
Attachment 3: err.png
err.png
  9166   Thu Sep 26 21:55:08 2013 ranaUpdateSUSProblems with ETMY Optical Lever

Not so fast! We need to plan ahead of time so that we don't have to repeat this ETMY layout another dozen times. Please don't make any changes yet to the OL layout.

Its not enough to change the optics if we don't retune the loop. Please do buy a couple of JDSU (and then we need to measure their intensity noise as you did before) and the 633 nm optics for the mode matching and then we can plan about the layout.

  9165   Thu Sep 26 11:00:51 2013 SteveUpdateSUSProblems with ETMY Optical Lever

We are out of JDSU-Uniphase 1103P heads. I'm ordering some right now. I'm planning to make some corrections on Rana's list tomorrow morning at ETMY.

  9164   Thu Sep 26 10:55:29 2013 SteveUpdateGeneralY-arm floor space for test

Liyuon will set up a  ~5 mW He/Ne laser for  waist measurement for LIGO oplev telescope.

This  will be between the beam tube and the CES wall. He will do his tests in the morning.

  9163   Thu Sep 26 01:49:28 2013 MasayukiSummaryGreen LockingFPMI noise caused by ARM locking

 

 Measurement with FPMI


i)By locking the FPMI with AS55Q and arms using POX,POY we measured  the OLTF on AS55Q, the response from BS actuation to error signal on AS55Q  for H_mich. The fitted,  measured OLTF and the residual function is in attachment1. I fitted two parameters and they are time-delay and the gain. The time delay is -275 usec. The time delay in three different control are almost same. The response from BS to AS55Q is in attachment 2.


With these two measuremets, I calclated the H_mich in FPMI. This H_mich should be different from simple MI because the cavity  refrectivity is different from the front mirror. Acrually it changed and the value was
Hmich = 4.4026e7

ii) I excited the ETMX and ETMY and measure the response from actuation to the error signal of MICH on AS55Q. The response is in attachment 3 and 4. from these result I calculated the H_L-l by using the formula as I mentioned. The value was
H_Lx-l = 175.7650 (XARM)
H_Ly-l = 169.8451 (YARM)

iii) I measured the error signal of MICH and XARM and YARM and with measured H_L-l, I estimated the FPMI noise caused by ARM locking. You can see in the higher frequency region than 10 Hz is dominated by noise caused by ARM control in-loop noises. 150 Hz and 220Hz are the UGF of each arms, so the two peaks are caused by arm control. You can see the small difference between FPMI noise and  noise from arms. There are two possibilities, one is that these measurement is not same time measurement so they should have small difference. and  other possibility is the error of the caliculation. But I think it doesn't look so bad estimation.

 

Next step

We will do same measurement with lock the arms the ALS system on tomorrow. Then we will check the PDH servo or other noise source and investigate the ALS system

Attachment 1: MICH_OLTF.pdf
MICH_OLTF.pdf
Attachment 2: BS-RS55Q.png
BS-RS55Q.png
Attachment 3: ETMX-RS55.png
ETMX-RS55.png
Attachment 4: ETMY-RS55.png
ETMY-RS55.png
Attachment 5: plot.pdf
plot.pdf
  9162   Wed Sep 25 23:59:29 2013 MasayukiSummaryGreen LockingFPMI noise caused by ARM locking

 

Measurement with ARMs

i) By locking the MICH with AS55Q signal I measured the actuator response of ITMX ITMY BS for calibration of each actuator. This measurement was done at the same time with elog#9158. The actuator response was
 
BS : 2.2347e-8 / f^2 [m/count]

ITMX: 5.0843e-9 /f^2 [m/count]

ITMY: 4.9677e-9 / f^2 [m/count]

 
 
ii)By locking the Arms for IR with POX,POY. I measured the OLTF and the response from ITM actuation to POX and POY signal. Attachment 1,2 are the plots of fitted OLTF , the measured OLTF, and residual function (model - measure)/model and the attachment 3,4 are the response of each arm. I fitted the three parameters. Those are the gain, time-delay and cavitypole. Each fitted parameter is
 
XARM ;
timedelay:-282.09 usec, cavity pole : 2872.0 Hz
YARM ;
timedelay:-283.84 usec, cavity pole : 2939.9 Hz
 
The cavity pole seems higher than privious measurement (In 2009). Actually the residual function are increase at the higher frequency region than 1kHz, so I guess the fitting is not so good.One possibility is that in the region where cavity pole effect increase we has not much data.
With fitted OLTF and actuator response I calibrated the H_xarm and H_yarm.
 
Hxarm : 2.9796 e11 [count / m]
Hyarm : 6.1394 e11 [count / m]
 
iii) After that I measured the response from ETM actuation to POX and POY signal to calibrate the ETM actuator. The response of each actuator is
 
ETMX:1.2040e-8 / f^2 [m/count]

ETMY:1.4262e-8 / f^2 [m/count]
 
iv) I calibrated the error signal with OLTF and Hxarm,Hyarm. The result is in Attachment 5

 In high frequency region there is the difference between xarm and yarm. These difference are already there in error signal. I'm not sure where these noise comes from. We will make measurement with Green PDH from tomorrow, so  we can also check with those measurement.

In other region the two noises are very close and also very similar to the plot of the seismic motion in the control room (attached on the front of TV screen).

Attachment 1: XARM_OLTF.pdf
XARM_OLTF.pdf
Attachment 2: YARM_OLTF.pdf
YARM_OLTF.pdf
Attachment 3: XARM_ITMXresponse.png
XARM_ITMXresponse.png
Attachment 4: YARM_ITMYresponse.png
YARM_ITMYresponse.png
Attachment 5: free_running.pdf
free_running.pdf
  9161   Wed Sep 25 23:15:11 2013 MasayukiSummaryGreen LockingFPMI noise caused by ARM locking
I measured some error signal, OLTFs and responses for FPMI noise estimation. Especially we are interested in the noise from in-loop noise of ALS Green PDH control. The strategy and
 
1) Purpose
 Estimation of the FPMI phase shift noise caused by in-loop noise of Green PDH control. 
 
 
2) What we should figure out
 For that estimation we have to figure out the transfer function from the cavity length change to the phase shift which is measured by MICH.
 
 
3) Strategy
 I attached the block diagram of  our interferometer. Our goal is to find the transfer function H_L-l and to calibrate the out of loop noise of interferometer with that TF and error signal of the PDH control.
 H,A and F mean the sensitivity, actuator response and servo filter for each control loop. L_xarm is the disturbance of the cavity length and l- is the differencial motion of the interferometer
fpmibd.pdf
We can get this H_L-l from measurement of the response from calibrated ETM actuation to the MICH error signal. You can get the formula for calculating H_L-l with simple calculation and that is
 
             1 + G_mich       1 + G_xarm      V_mi  
H_L-l = ---------------  -----------------  ------------
              H_mich             A_etmx         V_excetm
 
 
where the each G is OLTF and V_mi/Vexcetm is the response from the ETM actuation to the MICH error signal.
And then  the FPMI noise in the unit of meter / rHz is
 
                           H_L-l
N_fpmi = l_dis + ------------ Vx
                          H_mich
               
This second term is what we are interested in.
 
To estimate these noises
i) We can calibrate the actuators of  ITMX, ITMY and BS with using the MICH as sensor. So we can calibrate the arm error signals by  excitation of arm length using ITMs actuator.
ii) If we know the TFs of arms, we can calibrate the ETMX and ETMY actuators.
iii) We should know the response from ETMX or ETMY actuating to error signal of mich.
iv) Also we should calibrate the error signal of MICH in FPMI locking(H_mich). We can do that by exciting the BS.
 

Then we can estimate the noises.

 
In next entry, I will write about measurement.

 

  9160   Wed Sep 25 19:34:51 2013 ranaUpdateSUSProblems with ETMY Optical Lever

I went down to investigate the issue with the extra noise that I found in the ETMY optical lever yesterday. There were several problems with the optical layout down there - I'm not sure if I remember them all now.

  1. Beam reflected from OL QPD not dumped.
  2. OL QPD set normal to the steering mirror so that the back reflection goes into the vacuum chamber.
  3. HeNe laser mount only dogged with 2 dogs. Needs 3. Looks like some said "Aw, that's not goin' nowhere. Let's just leave that there pard!"
  4. First lens downstream of the laser had 2 screws and washers, but neither was even finger tight! They were loose by more than 1 full turn.
  5. Second lens was clipping. Beam was so far off center that this lens was being used to steer the beam by a few inches on the QPD.
  6. Extra reflections from ingoing beam (I don't know which surfaces) randomly landing on green & red optics.
  7. Lenses for the HeNe mode matching are coated for 1064 nm. HeNe is 633 nm, so these lenses must be replaced to reduce the reflections.

The main noise issue, however, appears to be not a layout issue at all. Instead its that the laser intensity noise has gone through the roof. See attached spectra of the quadrants (this is the way to diagnose this issue).

I'll ask Steve to either heal this laser or swap it out tomorrow. After that's resolved we'll need another round of layout fixing. I've done a couple of hours today, but if we want a less useless and noisy servo we'll have to do better.

NOTE: by looking at the OL quadrants, I've found a noisy laser, but this still doesn't explain the excess noise in the ETMX. That was the one that has a noisier error signal, not ETMY. By the coherence in the DTT, you can see that the ETMY OL is correctly subtracting and normalizing out the intensity noise of the laser. Seems like the ETMX electronics might be the culprit down there.

Attachment 1: ETMY-BadHeNe.pdf
ETMY-BadHeNe.pdf
  9159   Wed Sep 25 17:07:08 2013 ranaFrogsTreasureFree Green Mango Juice in fridge

aam_pana_recipe.jpg

its an acquired taste, but its a must since we're sending an interferometer to India

  9158   Wed Sep 25 08:11:01 2013 MasayukiSummaryLSCLSC calibration screen

 

  The real time calibration system is not correct in high frequency.

The attachment are the plot of two free running noise. Blue curve is the plot of noise calibrated with OLTF. Green one is the just fft analysed signal of the real time calibration system output. You can see the ripple in high frequency region in green curve. That is because the anti-aliasing filter and digital anti-aliasing filter. I assume the sensitivity of MI as constant but Rana mentioned that we should take these filters into account.

modeled OLTF and sensitivity H
 I put the AA filter and DAA filter effects into matlab calibration script. The attachment 2 is the modeled sensitivity of the MICH. You can find each filter properties in  elog#8555 (analog AA filter) and in elog#3961. I estimate the H gain by measuring the fringe. The attachment 3 is the plot of fringe and I averaged with green points. The actual number is 3.48e7 count/m

attachment 2: the sensitivitiy of MICH

Screenshot.png

 attachment 3: fringe of the MICH

fringe.pdf

I modeled OLTF with this H and the fitted into the measurement data. That is in attachment 4. In this OLTF I also included the DAI filter and AI filter, and ' sample and hold circuit' of DAC TF . These are  mentioned in two references. Additionally I added the time delay 309.6 us.  Yuta mentioned that in C1SUS has 125us time delay. In MICH control we have also C1LSC , so I think this time delay is reasonable. I compensated the error signal with these OLTF and MICH sensitivity.

attachment 4: OLTF of the MICH control

Screenshot-1.png

You can see that the ripple is gone in blue curve and after 5 kHz the curve is flat.

Next step

I'm trying to put the inverted AA filter and DAA filter in C1CAL_INCV servo. But the ploblem is the difference of sampling frequency, so I couldn't fix yet. One possibility is putting approximated filter. I hope I will find some good way to design these filters.

 

 

Other thing

I esitimated the FPMI noise propagated from the residual noise of IR PDH control of both ARMS. I will summarize and write these staff in this afternoon.

Attachment 1: plot.pdf
plot.pdf
  9157   Tue Sep 24 22:19:57 2013 ManasaUpdateGeneralFPMI locked

 [Masayuki, Manasa]

We locked FPMI and measured the FPMI noise (power spectrum of error signal - MICH_IN1) which will be calibrated.

The arms were locked using POX11 and POY11. The sign of MICH gain was changed to lock FPMI (from -30 to +30). 

  9156   Tue Sep 24 20:43:45 2013 masayukiSummarySUSoptical levers centering

I centered optical levers of ITMX,BS,ETMY. I also change the position of optical levers of ITMX, ETMY, ITMY, BS on Friday night(9/21), of ITMX, ETMY, BS on Monday night. Both are around 6:00 ~ 7:00.But centering on Monday was totally wrong, because I centered with not good IFO alignment.

The attachment is the 5 days trend of the opt lev of ITMX. First gap is alignment on Friday and Second gap is the alignment on Monday. Now I centered after  locking the FPMI.

The attachment 2 is the last 6 hours data.  The gap on 9/25 00:00 and 1:30(UTC)  is because the alignment of the cavity and the last gap is because of  centering of the optical lever.

Attachment 1: Screenshot-Untitled_Window.png
Screenshot-Untitled_Window.png
Attachment 2: Screenshot-Untitled_Window-1.png
Screenshot-Untitled_Window-1.png
  9155   Tue Sep 24 10:55:45 2013 ranaUpdatePSLPMC re-aligned

After relocking the PMC at a good voltage, Steve and I re-aligned the beam into the PMC by walking the last two steering mirrors. After maximizing the power, we also aligned the reflected beam by maximizing the PMC_REFL_DC with the unlocked beam.

Transmission is back to 0.84 V. We need Valera mode matching maintenance to get higher I guess. Maybe we can get a little toaster to keep the PMC PZT more in the middle of its range?

Attachment 1: psl-trend.png
psl-trend.png
  9154   Sun Sep 22 23:04:52 2013 ranaUpdatePEMGuralp needs recentering

 After seeing all of these spikes in the BLRMS at high frequency for awhile, I power cycled the Guralp interface box (@ 10:21 PM) to see if it would randomly recenter in a different place and stop glitching.

It did - needs to be better centered (using the paddle). Plot shows how the Z channel gets better after power cycle.

Attachment 1: seis.pdf
seis.pdf
  9153   Sun Sep 22 22:54:28 2013 ranaUpdateIOOmode cleaner not locking

Having trouble again, starting around 1 hour ago. No one in the VEA. Adjusted the offset -seems to be OK again.

  9152   Sun Sep 22 22:05:10 2013 ranaUpdateSUSoplev XY-plots reflect new calibration

The ETMX oplev signal looks kind of dead compared to the ETMY. It has no features in the spectra and the SUM is pretty low.

I noticed that the cal fields are still set to 1. To get it close to something reasonable, I calibrated it vs. the SUSPIT and SUSYAW values by giving it a step in angle and using 'tdsavg' plus some arithmetic.

OLPIT =  45 urads/ count

OLYAW = 85 urads / count

These are very rough. I don't even know what the accuracy is on the OSEM based calibration, so this ought to be redone in the way that Jenne and Gabriele did before.

The attached image shows the situation after "calibration" of ETMX. This OL system needs some noise investigation.

Attachment 1: noise.png
noise.png
  9151   Sun Sep 22 21:28:53 2013 ranaUpdateSUSset OL T RAMP values (they are not visible on the OL screens)

controls@rosalba:/opt/rtcds/caltech/c1/scripts/SUS 0$ ./setOLtramps
Old : C1:SUS-ETMX_OLPIT_TRAMP        0
New : C1:SUS-ETMX_OLPIT_TRAMP        2
Old : C1:SUS-ETMX_OLYAW_TRAMP        0
New : C1:SUS-ETMX_OLYAW_TRAMP        2
Old : C1:SUS-ETMY_OLPIT_TRAMP        2
New : C1:SUS-ETMY_OLPIT_TRAMP        2
Old : C1:SUS-ETMY_OLYAW_TRAMP        2
New : C1:SUS-ETMY_OLYAW_TRAMP        2
Old : C1:SUS-ITMX_OLPIT_TRAMP        0
New : C1:SUS-ITMX_OLPIT_TRAMP        2
Old : C1:SUS-ITMX_OLYAW_TRAMP        0
New : C1:SUS-ITMX_OLYAW_TRAMP        2
Old : C1:SUS-ITMY_OLPIT_TRAMP        0
New : C1:SUS-ITMY_OLPIT_TRAMP        2
Old : C1:SUS-ITMY_OLYAW_TRAMP        0
New : C1:SUS-ITMY_OLYAW_TRAMP        2
Old : C1:SUS-BS_OLPIT_TRAMP          0
New : C1:SUS-BS_OLPIT_TRAMP          2
Old : C1:SUS-BS_OLYAW_TRAMP          0
New : C1:SUS-BS_OLYAW_TRAMP          2
Old : C1:SUS-PRM_OLPIT_TRAMP         0
New : C1:SUS-PRM_OLPIT_TRAMP         2
Old : C1:SUS-PRM_OLYAW_TRAMP         0
New : C1:SUS-PRM_OLYAW_TRAMP         2
Old : C1:SUS-SRM_OLPIT_TRAMP         0
New : C1:SUS-SRM_OLPIT_TRAMP         2
Old : C1:SUS-SRM_OLYAW_TRAMP         0
New : C1:SUS-SRM_OLYAW_TRAMP         2
 
Done setting TRAMPs

  9150   Sun Sep 22 21:03:15 2013 ranaConfigurationSUSTuned bounce and roll mode of ETMY suspension

 Today I noticed that there was a lot of noise at the Bounce and Roll eigenfrequencies for ETMY. I found that the bandstop filter were set at completely the wrong frequencies, so I've remade them.

The filters were last tuned by Leo in May of 2011. Even so, he left the frequencies at the frequencies of the old MOS suspensions which had f_bounce ~ 12 Hz.

 The FOTON plot shows the OLD ones versus the NEW ones. The DTT spectra shows the oplev error signals in the usual state. I have also copied these over to the SUSPOS,PIT,YAW, and SIDE filter banks and turned them all ON.

I also turned OFF and deleted the 3 Hz RG filter that was there. There's no such peak in the error signal and even if one wanted to compensate for the stack mode, it should be a low Q filter, not this monster.

Attachment 1: etmy-error.png
etmy-error.png
Attachment 2: notches.pdf
notches.pdf
  9149   Fri Sep 20 22:49:10 2013 MasayukiUpdateLSCLSC calibration screen

Quote:

I update the LSC calibration screen. This screen is for real time calibration of each DOF with using error signal and control signal. The formula of the calibration is

x_dis = V_err/H + A V_fb

,where x_dis is the disturbance without surpression, V_err and V_fb are error signal and control signal, H is the transfer function from the displacement to output and A is the efficiency of the actuator.

I will put the filter of 1/H into the CINV filter bank and actuator efficiency into the A filter bank.

 

 I fixed the filter of the MICH real-time calibration. You can find C1CAL screen from the LSC menu 'calibration' of sitemap.

*Filter explanation

    C1CAL_MICH_CINV : the servo to convert  the error signal to displacement.

Sen_MICH :

the inverse of the transfer function from the distance to the error signal, which has the unit of count/m. In the formula this filter is represented by 1/H.

I assume this H is independent of frequency and time, and I calculated by the amplitude of the fringe of error signal. But it may change every day by drift of laser intensity and so on.  So we should follow the actual H somehow.  The temporary value of H is 3.76*10^7 count/m .

    C1CAL_MICH_A : the servo to convert the feedback signal to displacement. In formula This transfer function is represented by A

 SUS_BS;

the transfer function of the suspension of the BS. This is modeled from the measurement in elog#9127. The resonant frequency is 1.029 Hz and Q is 12.25.

 Res_A :

the response of the actuator on BS_SUS, which has the unit of m/count. The value is 1.99*10^-8 m/count. This value is measured in the measurement in elog#9121.

     C1CAL_MICH_W : the servo to handle the calibrated signal.

  m->um ;

the filter to convert the unit of signal from m to um. When this filter is on, the output is written in unit of um.

*Measurement

 I measured the power spectrum of the calibrated free running noise. The measured port was C!CAL_MICH_W_OUT. The result is in attachment 1. Also in this figure there are the plots of the Verr/H and Vfb*A.

 In low frequency region, where control loop suppresses the disturbance, you can see that the displacement is equal to the displacement of actuation (I'm not sure what happens at the point of 0.03Hz), and in high frequency region, where control loop doesn't work, the displacement is equal to the value of the Verr divided by MICH sensitivity. Also this result is similar to the my calibration result.elog#9131

Attachment 1: plot.pdf
plot.pdf
  9148   Fri Sep 20 20:27:18 2013 ranaUpdateIOOmode cleaner not locking

 I used our procedure from this entry to set the IMC board offset as well as the FSS board offset.

I found this afternoon that the MC was having trouble locking: the PC path was railing as soon as the boost was engaged. Could be that there's some misalignment on the PSL which has led to some RAM having to be canceled by this new offset. Let's see if its stable for awhile.

  9147   Fri Sep 20 20:14:52 2013 ranaSummaryGeneral/home/cds nearly full

Quote:

One of the reasons that our disk is getting full is due to the scripts_archive directory. A backup script runs on op340m and makes a tar.bz2 file of the scripts directory and puts it in scripts_archive every morning at 6 AM.

On Oct 7, 2011, Koji fixed this script to point at our new scripts directory instead of the old /cvs/cds/caltech/scripts directory. Since then, however, no one has fixed the exclude file to NOT back up the junk that's in that directory. Its a 1.6 GB directory so its full of it.

I've deleted a bunch of junk from the scripts directory: this directory is for scripts, not for your personal home movies or junk data files. Put those in your USER directory. Put temporary data files in /tmp/. I've also added a few more patterns to the exclude file so that less .mpg, .png, .pdf, .dat, etc get stored every day. The new daily .tar.bz2 file wil be ~25 MB instead of 770 MB.

(also fixed the backup script to use 'env' to setup the perl environment and removed the hard-coded path to tar)

 OUr disk was getting full again. Turned out my "fix" to 25 MB was only a fix to 250 MB. Since we were getting disk full warnings on our Ubuntu workstations, I deleted some COMSOL.dmg files from users/zach/ and then started deleting every other tarball from the scripts_archive directory. ~221 GB are now free. Still need to fix the exclude file for scripts better.

  9146   Fri Sep 20 16:37:30 2013 SteveUpdateVACvertex crane folding arm is back on

Quote:

 The folding beam removed as shown. Two man supporting it while I hammering it out. Pin was dry and it gulled into supporting hinges.

The rotating hinge will be machined and bushing will be added with Zerk fitting or similar. This will allow lubrication in the future.

 see elog #9111

Atm1,  The folding arm is back on with 0.1" misalignment at no load in the trolly's way. The other side of the I beam is 0.02" higher than the main beam.. New bushing and pin were greased up  with Krytox before installation.

The axial Zerk 1/8" pipe in the pin upper end can not take any fitting. There is no room. It is taped off.

This gap comes down to ~ 1/16" at fully extended arm with 225 lbs load at the end of it.

The present plan is to grind down the  the misalignment of 0.1"  for a slow-loaded trolly.

Steve Baker of Konacranes will be back to grind down this ridge and load test at 500 lbs on Tuesday,  OCT  1, 2013

 

 

 

Attachment 1: bushingBottom.jpg
bushingBottom.jpg
Attachment 2: foldingI.jpg
foldingI.jpg
Attachment 3: noLoadGap.jpg
noLoadGap.jpg
Attachment 4: trollyway.jpg
trollyway.jpg
Attachment 5: trollySee.jpg
trollySee.jpg
ELOG V3.1.3-